After a genetics experiment on 63 pea plants, the number of plants having certain 23 were tall:; characteristics was tallied, with the results given to the right. Answer parts (a) through (c). 27 had green peas; 44 had smooth peas; 8 were tall and had green peas; (a) Find the number of plants that were tall and had smooth peas. 18 had green peas and smooth peas 3 had all three characteristics; 6 had none of the characteristics. plants(a) Find the number of plants that were tall and had smooth peas. (b) How many plants were tall and had peas that were neither smooth nor green? (c) How many plants were not tall but had peas that were smooth and green?
Answer #1

Answer=

had all three characteristics = 3
tall and had green peas = 8- 3= 5
tall and had smooth peas = 18 - 3= 15
had green peas and smooth peas = 18 - 3= 15
tall only = 23-3- 5- 15= 0
had green peas only = 27- 3- 5-15= 4
had smooth peas only = 44- 3- 15- 15= 11
none of the characteristics = 6

so the total # of plants counted =3+5+15+15+4+11+6=59

(a)No. of plants that are tall and smooth peas=15

(b)plant which are tall neither smooth nor green=0

(c)plants green and smooth=15